OpenStudy (anonymous):

integrate x^3/(x^4+1)....seems like a partial fraction kind of problem but I can't find how to solve

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ y^3 }{ y^4 + 1}\]

hartnn (hartnn):

there's a lot easier way try to substitute u= y^4+1 what do u get du as ?

OpenStudy (anonymous):

I got half a page of the wrong answer.....I'll try the usub again brb

hartnn (hartnn):

tell me what u get ?i may be able to point out the error..

OpenStudy (anonymous):

wolf says my answer should be 1/4 ln(x^4 +1) .... that wrong answer is a crumpled ball on the other side of the room i'm starting over

hartnn (hartnn):

good, so 1st step is to find du

OpenStudy (anonymous):

ah dy = 1/(4x^3) so the x^3's cancel thats where I screwed up --- I didn't like the x^3 in the numerator and panicked

OpenStudy (anonymous):

that is why an algorithm based on if you can't use usub try parts and then try fractions....me not being able to usub is a dodgy criteria

hartnn (hartnn):

so you could solve this completely now right? if you get stuck again, you can ask...

OpenStudy (anonymous):

the integration is done....the differential solution is a different problem but I'll start a new Q if I can't figure out how that ='s sin(y) or maybe that is the answer....can't e both sides...does e^sin(x) mean anything...i'll work on it thanks a ton

hartnn (hartnn):

ok welcome ^_^